OpenStudy (anonymous):

Part 2: Using these similarities and differences, how would you transform f(x) = 2 sin(2x - π) + 3 into a cosine function in the form f(x) = a cos(bx - c) + d?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

any thoughts?

OpenStudy (anonymous):

Part 1: Using complete sentences, compare the key features and graphs of sine and cosine. What are their similarities and differences? The graph of cosine and sine are pretty much the same except that the graph of cosine is shifted by pi / 2 behind the sine graph. I did that for part 1

OpenStudy (phi):

for part 2, both sin and cos have the same shapes (just shifted), so the numbers on the outside should match up. i.e. the a and d

OpenStudy (anonymous):

cosine is just shifted from sine? im guessing we will use pi/2

OpenStudy (anonymous):

and we need to combine the function somehow but im not quite sure what to plug in

OpenStudy (phi):

first a=2 and d is 3

OpenStudy (phi):

next we know sin(x) = cos(x-π/2) to make it clearer (maybe). call 2x - π = A sin(A) = cos(A - π/2) now replace A with 2x-π sin(2x - π) = cos(2x-π - π/2)

OpenStudy (anonymous):

ok i understand all of that, what would be next? simplify that?

OpenStudy (phi):

I would simplify -pi - pi/2 to be -2pi/2 - pi/2 = -3pi/2 so you get sin(2x - π) = cos(2x-3π/2 )

OpenStudy (anonymous):

ok i understand how you simplified

OpenStudy (phi):

now match cos(bx - c) with cos(2x-3π/2 )

OpenStudy (phi):

b is 2 and c is 3pi/2

OpenStudy (anonymous):

cos(2x - 3pi /2) they're the same?

OpenStudy (phi):

?

OpenStudy (phi):

the final result is f(x) = 2 sin(2x - π) + 3 becomes f(x) = 2 cos(2x - 3π/2) + 3?

OpenStudy (anonymous):

this is confusing do i use the full formula then? f(x) = a cos(bx - c) + d 2cos(2x - 3pi/2) +3?

OpenStudy (anonymous):

okay i'll try to sum all of this up, was part 1 right? up near the top

OpenStudy (phi):

they want to rewrite 2 sin(2x - π) + 3 using cosine instead of sine we know sin(A)= cos(A - π/2) so we "snip out" sin(2x - π) and put in cos(2x - π - π/2) instead: 2 cos(2x - π - π/2)+ 3 and simplify the "inside" part to get 2 cos(2x - 3π/2)+ 3

OpenStudy (anonymous):

wow thanks once again, i couldn't do this without you haha. i have a couple more if you wouldn't mind going over

OpenStudy (phi):

what I wrote is confusing. better is the cosine is the sin curve shifted to the left pi/2 radians

OpenStudy (phi):

I have to sign off for now.

OpenStudy (anonymous):

okay thanks for the help!